OpenStudy (anonymous):

PLEASE HELP !!!! DUE IN 5 MINUTES . GIVING MEDALS TO ALL CONTRIBUTORS !! Solve. 5 = –(–z + 3)

OpenStudy (anonymous):

5= +z - 3 z= 5+3 z= 8

OpenStudy (twopointinfinity):

You can only give 1 medal.